OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (anonymous):

x^2=-1 if u define an imaginary solution as no solution.

OpenStudy (anonymous):

or, 0*x=1

OpenStudy (anonymous):

I think 0*x=1 is better, because if u try to solve it ull get x=1/0

OpenStudy (anonymous):

and you can't divide by 0

OpenStudy (accessdenied):

 at least, thats what im thinking i imagine we're just assuming we're only looking in the Reals